### v1

# A Model-Based Derivative-Free Optimization Algorithm for Partially Separable Problems

Abstract.
We propose UPOQA, a derivative-free optimization algorithm for partially separable unconstrained problems, leveraging quadratic interpolation and a structured trust-region framework. By decomposing the objective into element functions, UPOQA constructs underdetermined element models and solves subproblems efficiently via a modified projected gradient method. Innovations include an approximate projection operator for structured trust regions, improved management of elemental radii and models, a starting point search mechanism, and support for hybrid black-white-box optimization, etc. Numerical experiments on 85 CUTEst problems demonstrate that UPOQA can significantly reduce the number of function evaluations. To quantify the impact of exploiting partial separability, we introduce the speed-up profile to further evaluate the acceleration effect. Results show that the acceleration benefits of UPOQA are marginal in low-precision scenarios but become more pronounced in high-precision scenarios. Applications to quantum variational problems further validate its practical utility.
Partially separable problems, derivative-free optimization, trust region methods, interpolation model
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Mathematics of computing Solversccs: Mathematics of computing Nonconvex optimization
1. Introduction
Derivative-free optimization (DFO) (Conn et al., 2009) methods, also known as black-box optimization (BBO) or zero-order optimization methods, are designed to find the minimum or maximum of an objective function when its derivatives are not available. For complex problems, the objective function may not be expressed as a closed-form analytical expression but rather as a black box that only allows function evaluations. Examples include cases where function values are obtained from simulation or experimental observations, or from the output of software packages with unknown internal implementations. Such objective functions often incur high evaluation costs, yield noisy values, or both.
DFO algorithms can be broadly categorized into five classes. The first class comprises direct search and pattern search methods (Lewis et al., 2000; Lewis and Torczon, 1999), which select the next iteration point by comparing the relative merits of function values (Hock and Schittkowski, 1980). These methods typically employ specific geometric patterns for point selection. They impose minimal requirements on the properties of the objective function, are computationally simple, and less prone to local minima, but often exhibit slow convergence. The second class consists of line-search-based methods, with Powell’s method (Powell, 1964) being a representative example that performs line searches along conjugate direction sets. The third class includes heuristic search methods, which explore the solution space following specific heuristic rules to find approximate optima. Common algorithms in this category are simulated annealing (Kirkpatrick et al., 1983) and evolutionary algorithms (Runarsson and Yao, 2005; da Silva Santos et al., 2010), with CMA-ES being a notable example (Hansen, 2006). The fourth class is model-based derivative-free optimization (Karasözen, 2007; Conn and Toint, 1996; Ragonneau and Zhang, 2024; Cartis et al., 2019). These methods construct a surrogate model through interpolation or regression using known function values, approximating the objective function locally. The model is then optimized and updated within a trust-region framework to solve the original problem. Examples include Powell’s COBYLA (Powell, 1994) (based on first-order models) and BOBYQA (Powell, 2009), NEWUOA (Powell, 2006) (based on second-order models), etc. Surrogate models may also adopt other forms, such as radial basis functions (Oeuvray, 2005) and moving ridge functions (e.g., OMoRF (Gross and Parks, 2022)). The fifth class comprises finite-difference-based methods(Berahas et al., 2022), which approximate the true gradient using numerical differentiation and then employ gradient-based optimization algorithms.
In DFO, the primary challenge lies in the lack of structural knowledge about the problem. When the problem lacks smoothness, convexity, or is subject to significant noise, many DFO methods essentially sample the parameter space according to certain patterns. Such approaches are prone to the curse of dimensionality when dealing with large-scale problems. Even if smoothness is assumed and surrogate models are constructed to accelerate convergence, the computational cost of solving these problems may still be prohibitive (Ros and Hansen, 2008). To address this, many researchers have considered structured DFO problems, such as nonlinear least squares (Powell, 1965; Cartis et al., 2019; Zhang and Conn, 2012; Hough and Roberts, 2022), problems with sparse Hessian matrices (Colson and Toint, 2002, 2001; Ros and Hansen, 2008) and partially separable problems (Porcelli and Toint, 2022; Price and Toint, 2006; Bouzarkouna et al., 2011; Shahabuddin, 1996; Colson and Toint, 2005; Zhu et al., 2019). Exploiting these structures can significantly reduce the number of function evaluations required for convergence.
This paper focuses particularly on unconstrained partially separable problems (Griewank and Toint, 1982a), which take the form:
(1) $\min_{x\in{\mathbb{R}}^{n}}f(x)=\sum_{i=1}^{q}f_{i}(x),$
where $f_{1},\ldots,f_{q}$ are referred to as element functions (or simply elements). For each $i=1,\ldots,q$, there exists a subspace $\mathcal{N}_{i}\subset\mathbb{R}^{n}$ such that for any $w\in\mathcal{N}_{i}$ and $x\in\mathbb{R}^{n}$, the following holds:
$f_{i}(x+w)=f_{i}(x).$
In contrast to the elements, $f$ is called the overall function. Let $\mathcal{R}_{i}\triangleq\mathcal{N}_{i}^{\perp}$, and denote $n_{i}\triangleq\dim\mathcal{R}_{i}$ as the elemental dimension of $f_{i}$. Each $f_{i}$ is essentially a function defined on $\mathcal{R}_{i}$. The corresponding projection is denoted as $P_{\mathcal{R}_{i}}\colon\mathbb{R}^{n}\to\mathcal{R}_{i}$, ensuring that for any $x\in\mathbb{R}^{n}$, $f_{i}(x)\equiv f_{i}(P_{\mathcal{R}_{i}}(x))$. Without loss of generality, we assume $\operatorname{span}\left(\cup_{i=1}^{q}\mathcal{R}_{i}\right)=\mathbb{R}^{n}$, as otherwise, redundant variables would not affect the value of $f$.
Partially separable structures frequently arise in optimal control problems, the discretization of partial differential equations (PDEs) via methods like finite elements, and of other variational problems (Griewank and Toint, 1982b). In optimal control, problems often consist of multiple loosely coupled subsystems distributed across space and time. Similarly, problems stemming from domain decomposition techniques applied to PDE discretizations exhibit partial separability, where each variable primarily interacts with its neighboring regions. In quantum computing, many variational quantum problems also demonstrate such structures, with each subfunction represented by a quantum circuit (Bierman et al., 2022, 2023).
In practice, the more commonly used concept is coordinate partial separability, which further requires that each $f_{i}$ depends only on a subset of variables $\left\{x_{j}\mid j\in\mathcal{I}_{i}=\{j_{1},\ldots,j_{n_{i}}\}\right\}$, referred to as the elemental variables of $f_{i}$. In this case, $\mathcal{R}_{i}=\operatorname{span}(\{e_{j}\mid j\in\mathcal{I}_{i}\})$, and we denote $P_{\mathcal{R}_{i}}(x)$ as $x^{\mathcal{I}_{i}}$. Since $f_{i}(x)$ depends solely on $x^{\mathcal{I}_{i}}$, we will, for simplicity, no longer distinguish between the notations $f_{i}(x^{\mathcal{I}_{i}})$ and $f_{i}(x)$ in subsequent discussions. Due to its prevalence and algorithmic convenience, unless explicitly stated otherwise, all references to partial separability in this paper refer to coordinate partial separability. We further assume that each element function can be evaluated independently. This assumption aligns with the practical context of most problems, as when each $f_{i}$ represents the contribution of a subsystem, these contributions should naturally be computable separately.
The term partial in partial separability refers to the fact that the variables on which different element functions depend may overlap. The two extremes of this property are when all elements depend on completely different variables, such as $f(x)=\sum_{i=1}^{q}f_{i}(x_{i})$, and when all elements depend on all variables, i.e., $f(x)=\sum_{i=1}^{q}f_{i}(x)$. However, most real-world problems lie between these two cases (Toint, 1983; Ding and McKee, 2011; Li et al., 2013). In the LANCELOT package (Conn et al., 2010), the concept of partial separability is further generalized to group partial separability and used to define the standard input format (SIF) for test problems. Moreover, many common problem structures can be viewed as special cases of partially separable structures. Any twice continuously differentiable function with a sparse Hessian can be transformed into a partially separable form (Griewank and Toint, 1982a), and nonlinear least squares problems $\min_{x}\sum_{i=1}^{N}r_{i}(x)^{2}$ are naturally partially separable.
The most obvious advantage of partial separability is that, given a trial point $x\in{\mathbb{R}}^{n}$, one can always compute the value of any element function $f_{i}(x)$ individually without evaluating the full $f(x)$. Conversely, even if $f(x)$ must be computed, all element values $f_{1}(x),\ldots,f_{q}(x)$ can be obtained simultaneously, thereby significantly increasing the amount of information available. Various methods have been developed to exploit partial separability. When derivatives are available, a common approach is to employ partitioned Hessian updating schemes (Griewank and Toint, 1982b, 1984; Toint, 1986; Lescrenier, 1988; Cao and Yao, 2016) in quasi-Newton methods, along with techniques such as element merging and grouping (Conn et al., 1994; Kaya et al., 2019), element-wise preconditioning (Daydé et al., 1997), and parallelization (Lescrenier, 1988) to reduce the computational cost of Newton steps or conjugate gradients. Conn et al. proposed a trust-region framework (Conn et al., 1996) for solving partially separable problems, which was later extended by Shahabuddin (Shahabuddin, 1996) with three algorithm variants based on different methods for solving the trust-region subproblems. Another approach is incremental optimization (Bertsekas et al., 2011; Kaya et al., 2019), where only a small subset of elements is optimized in each iteration.
When derivatives are unavailable, existing approaches can be broadly categorized into bottom-up and top-down approaches, depending on whether they directly optimize the element functions or optimize the objective function. The bottom-up approach primarily leverages the independent evaluability of elements to reconstruct information about the objective function from element values at relatively low cost, thereby solving problem (1). A straightforward implementation evaluates elements at grid points in $\mathbb{R}^{n}$, e.g., $\{x=$ $\left.\sum_{i=1}^{n}c_{i}e_{i}\mid c_{i}\in\mathbb{Z}\right\}$, then assembles these evaluations into full function values. For instance, for ${f(x,y,z)}=f_{1}(x,y)+f_{2}(y,z)$, computing the values of $f_{i}$ at $\{0,1\}^{2}$ requires only 4 evaluations but yields all 8 values of $f$ over $\{0,1\}^{3}$. If elements are independent, evaluating each element $k$ times produces $k^{q}$ objective values from just $kq$ evaluations. Price and Toint (Price and Toint, 2006) exploited this to develop a mesh-refinement-based direct search method that scales efficiently to thousands of dimensions. Another strategy is line-search-based, such as Porcelli and Toint’s BFO method (Porcelli and Toint, 2022).
An alternative approach is top-down, where the algorithm primarily extracts information from complete evaluations of $f$ to guide the optimization. Model-based methods often fall into this category, as the subproblems formed by surrogate models typically generate points without a clear geometric pattern, making it unlikely to reconstruct the objective value by computing only a few element values. Relevant work includes Bouzarkouna et al.’s p-sep lmm-CMA method (Bouzarkouna et al., 2011) and Conn and Toint’s PSDFO method (Colson and Toint, 2005). The p-sep lmm-CMA method is a variant of lmm-CMA (Kern et al., 2006), where a meta-model (a second-order polynomial regression model) is built for each element function to guide the ranking of data points in lmm-CMA. Meanwhile, PSDFO is an interpolation-based trust-region method where each element $f_{i}$ is approximated by a quadratic interpolation model $m_{i}$. In both methods, the primary purpose of modeling elements is to construct an overall model
(2) $m(x)=\sum_{i=1}^{q}m_{i}(x),$
which is then used as a single surrogate model in the underlying algorithm for general optimization. Consequently, the information available to the algorithm mainly stems from feedback on the full function $f$ or the overall model $m$. The benefits of the top-down approach are relatively indirect. Compared to using a single surrogate model, employing multiple element models primarily reduces the total modeling cost, since evaluating $f$ once yields all values $f_{1},\ldots,f_{q}$, enabling simultaneous updates to all models in a single iteration.
Generally speaking, model-based methods exhibit superior computational efficiency compared to direct search and pattern search approaches. In pursuit of efficiency, we adopt a top-down strategy and develop a trust-region algorithm that employs quadratic interpolation models to approximate element functions. The algorithm is named Unconstrained Partially-separable Optimization by Quadratic Approximation (UPOQA). The key features of UPOQA include:
  1. (1)
    For interpolation modeling, UPOQA utilizes the underdetermined quadratic model based on Powell’s derivative-free symmetric Broyden update (Powell, 2004a), incorporating Powell’s techniques (Powell, 2004b, 2008) to enable efficient model construction and maintenance with low algebraic complexity;
  2. (2)
    For trust-region management, UPOQA assigns independent trust-region radii to each element, implements a modified projected gradient method for solving trust-region subproblems within structured trust regions, improves upon the existing radius adjustment strategy (Shahabuddin, 1996), and introduces a selective update mechanism for element models based on obtained interpolation points;
  3. (3)
    In terms of accessibility, UPOQA provides a ready-to-use Python implementation with various features to enhance flexibility and robustness, including hybrid black-white-box optimization and restart mechanism. The open-source package is released on GitHub under the BSD 3-Clause License.
The rest of the paper is organized as follows. Section 2 presents the overall framework of UPOQA, including the designed approximate projection operator for structured trust-region problems, the management of trust-region radius, the criteria for selective element model updates, starting point search mechanism, and optional features (restart and hybrid optimization). Section 3 conducts comprehensive numerical experiments on CUTEst test problems extracted via the S2MPJ tool (Gratton and Toint, 2024) and test cases from the quantum variational problem (17) to validate the algorithm’s effectiveness. Section 4 concludes the work and discusses potential future research directions.
2. The UPOQA Algorithm
The UPOQA algorithm is based on Powell’s trust-region framework (Powell, 2006, 2009; Ragonneau and Zhang, 2024) for solving problem (1). For each element $f_{i}$, the algorithm maintains an interpolation set $\mathcal{Y}_{i}$ of size at least $O(n_{i})$, along with an underdetermined quadratic interpolation model (Powell, 2004a) constructed from these points:
(3) $m_{k,i}(x_{k}^{\mathcal{I}_{i}}+s)=c_{k,i}+g_{k,i}^{\top}s+\frac{1}{2}s^{\top}H_{k,i}s,\quad s\in{\mathbb{R}}^{n_{i}}.$
We refer to $m_{k,1},\ldots,m_{k,q}$ as element models, where $c_{k,i}\in{\mathbb{R}}$, $g_{k,i}\in{\mathbb{R}}^{n_{i}}$, and $H_{k,i}\in{\mathbb{R}}^{n_{i}\times n_{i}}$ is a symmetric matrix. The model $m_{k,i}$ is expected to sufficiently approximate the objective function within a trust region of radius $\Delta_{k,i}$:
$\mathcal{B}_{i}(\Delta_{k,i})=\left\{x_{k}^{\mathcal{I}_{i}}+s\mid s\in{\mathbb{R}}^{n_{i}},\ \|s\|_{2}\leq\Delta_{k,i}\right\}.$
Moreover, each model is equipped with its own trust-region radius $\{\Delta_{k,i}\}_{i=1}^{q}$ to account for their varying reliability due to different accuracies.
During the execution of UPOQA, as the interpolation sets are updated, the element models are also updated by the derivative-free symmetric Broyden update (Powell, 2004a), which has been widely employed by Powell in algorithms such as NEWUOA (Powell, 2006) and BOBYQA (Powell, 2009). It allows the algorithm to use only $O(n_{i})$ interpolation points per model while updating the model with $O(n_{i}^{2})$ algebraic complexity (Powell, 2004b, 2008). This surrogate model is widely recognized as an effective second-order approach in derivative-free methods based on interpolation models and trust regions (Conn et al., 2009).
Our work is tailored for partially separable problems. When the number of elements exceeds one, two key issues must be addressed. First, how to solve the resulting structured trust-region subproblem:
$\min\limits_{s\in{\mathbb{R}}^{n}}$ $m_{k}(x_{k}+s),$
s.t. $\|s^{\mathcal{I}_{i}}\|_{2}\leq\Delta_{k,i},\quad i=1,2,\ldots,q,$
where $m_{k}\triangleq\sum_{i=1}^{q}m_{k,i}$ is the overall model assembled from the element models $\{m_{k,i}\}_{i=1}^{q}$. Second, how to evaluate the contribution of each element model to the reduction of the objective value in each iteration and accordingly adjust $\{\Delta_{k,i}\}_{i=1}^{q}$. For the first issue, we employ a modified projected gradient method based on an approximate projection operator, which will be detailed in Section 2.1. For the second issue, we directly adopt the combined separation criterion proposed by Shahabuddin (Shahabuddin, 1996, Section 2.4), with minor modifications in corner cases.
Next, we briefly outline the general workflow of the UPOQA algorithm for solving problem (1). At initialization, the algorithm first constructs the surrogate model $m_{0,i}$ and the interpolation set $\mathcal{Y}_{0,i}$ for each element $f_{i}$. Then, using all interpolation points from $\{\mathcal{Y}_{0,i}\}_{i=1}^{q}$, it seeks a better starting point $x_{0}$ via the method described in Section 2.4. If successful, the objective value $f(x_{0})$ will be lower than that of the user-provided point $x_{\text{start}}$. In each iteration, UPOQA first solves the trust-region subproblem (2) using the modified projected gradient method from Section 2.1, obtaining a trial step $s_{k}$. The algorithm then decides whether to accept this step based on the magnitude of $\|s_{k}\|_{2}$.
For $s_{k}$ with large norms, the algorithm first computes a candidate point $\hat{x}_{k}=x_{k}+s_{k}$ and updates all models and interpolation point sets based on $\hat{x}_{k}$. When updating the model $m_{k,i}$ and the interpolation set $\mathcal{Y}_{k,i}$, a new interpolation point $\hat{x}_{k}^{\mathcal{I}_{i}}$ is added to $\mathcal{Y}_{k,i}$. However, this procedure is performed only if $\hat{x}_{k}^{\mathcal{I}_{i}}$ does not compromise the poisedness of the interpolation set $\mathcal{Y}_{k,i}$. For updating the trust-region radii $\left\{\Delta_{k,i}\right\}_{i=1}^{q}$, we employ the combined separation criterion proposed by Shahabuddin (Shahabuddin, 1996, Section 2.4). If the objective exhibits insufficient reduction at $\hat{x}_{k}$, the algorithm further checks whether a geometry-improving step should be taken. If the conditions for a geometry-improving step are not met, the trust-region resolution $\rho_{k}$ is reduced. The value of $\rho_{k}$ serves as a lower bound for all trust-region radii $\{\Delta_{k,i}\}_{i=1}^{q}$, indicating the search precision of the algorithm at the current iteration. For $s_{k}$ with small norms, the algorithm will reduce the trust-region radii and perform geometry-improvement steps. If such small-step events occur consecutively multiple times or no geometry improvement steps are triggered, the algorithm also reduces $\rho_{k}$.
Algorithm 1 outlines the procedure of UPOQA. Steps unrelated to the partially separable structure closely resemble those in Powell-style model-based algorithms. For these aspects, we refer readers to the descriptions of these algorithms (Powell, 2006, 2008; Ragonneau, 2023; Ragonneau and Zhang, 2024) or the implementation of UPOQA for further details.
Input: Objective $f$, elements $\{f_{i}\}_{i=1}^{q}$, index sets $\{\mathcal{I}_{i}\}_{i=1}^{q}$, starting point $x_{\text{start}}$, initial trust-region resolution $\rho_{0}$
Output: Approximate solution to problem (1)
1 Initialize $\{\mathcal{Y}_{0,i}\}_{i=1}^{q}$ and $\{m_{0,i}\}_{i=1}^{q}$ centered at $x_{\text{start}}$, set $\boldsymbol{\Delta}_{0}=[\rho_{0},\ldots,\rho_{0}]^{\top}\in{\mathbb{R}}^{q}$;
2
3Compute an improved starting point $x_{0}$ using the method described in Section 2.4;
4
5for $k=0,1,\ldots$ do
6       $F^{GI}_{1},\ldots,F^{GI}_{q}\leftarrow\textbf{false}$;
7
8      Solve the trust-region subproblem
$\min\limits_{s\in{\mathbb{R}}^{n}}$ $\sum_{i=1}^{q}m_{k,i}(x_{k}+s),$
s.t. $\|s^{\mathcal{I}_{i}}\|_{2}\leq\Delta_{k,i},\quad\text{for }i=1,2,\ldots,q,$
to obtain the step $s_{k}$ using the method detailed in Section 2.1 and set $\hat{x}_{k}\leftarrow x_{k}+s_{k}$;
9
10      if $\max_{i}\left(\left\|s_{k}^{\mathcal{I}_{i}}\right\|_{2}\right)\leq\rho_{k}/2$ then
11             for $i=1,2,\ldots,q$ do
12                   $\Delta_{k,i}\leftarrow\max\left(\Delta_{k,i}/2,\rho_{k}\right)$;
13
14                  if the conditions for geometry improvement of model $m_{k,i}$ are met then
15                         $F^{GI}_{i}\leftarrow\textbf{true}$;
16
17
18            if entered this branch in multiple consecutive iterations $\mathbf{or}$ $F^{GI}_{1},,\ldots,F^{GI}_{q}$ are all false then
19                   $\rho_{k}\leftarrow 0.1\rho_{k}$;
20
21
22      else
23             Compute
$r_{k}\leftarrow\frac{f(\hat{x}_{k})-f(x_{k})}{m_{k}(\hat{x}_{k})-m_{k}(x_{k})},\quad r_{k,i}\leftarrow\frac{f_{i}(\hat{x}_{k})-f_{i}(x_{k})}{m_{k,i}(\hat{x}_{k})-m_{k,i}(x_{k})},\quad\text{for }i=1,2,\ldots,q;$
24            for $i=1,2,\ldots,q$ do
25                   if $\hat{x}_{k}^{\mathcal{I}_{i}}$ does not significantly deteriorate the poisedness of $\mathcal{Y}_{k,i}$ then
26                         Select an interpolation point $y_{k}^{\text{del},i}$ to remove from $\mathcal{Y}_{k,i}$;
27                         Set $\mathcal{Y}_{k,i}\leftarrow\mathcal{Y}_{k,i}\backslash\left\{y_{k}^{\text{del},i}\right\}\cup{\hat{x}_{k}^{\mathcal{I}_{i}}}$ and update the element model $m_{k,i}$;
28
29                  Adjust the trust-region radius $\Delta_{k,i}$ using the method detailed in Section 2.2;
30
31            if $r_{k}\leq 0.1$ then
32                   for $i=1,2,\ldots,q$ do
33                         if the conditions for geometry improvement of model $m_{k,i}$ are met then
34                               $F^{GI}_{i}\leftarrow\textbf{true}$;
35
36                  if $F^{GI}_{1},,\ldots,F^{GI}_{q}$ are all false then
37                         $\rho_{k}\leftarrow 0.1\rho_{k}$;
38
39
40
41      Set $x_{k+1}$ as the point with the smaller function value between $x_{k}$ and $\hat{x}_{k}$;
42       Perform geometry-improving steps for all element models and interpolation sets where $F^{GI}_{i}=\textbf{true}$;
43
Algorithm 1 The UPOQA Algorithm
2.1. Structured Trust Region
UPOQA assigns an independent trust-region radius $\Delta_{k,i}$ to each element model $m_{k,i}$. Consequently, the feasible region of the trust-region subproblem is no longer a ball but rather an $n$-dimensional cylinder set defined by the projections $\left\{P_{\mathcal{R}_{i}}\colon{\mathbb{R}}^{n}\to\mathcal{R}_{i}\right\}_{i=1}^{q}$ and the regions $\left\{\mathcal{B}_{i}(\Delta_{k,i})\subset\mathcal{R}_{i}\right\}_{i=1}^{q}$:
(5) $\mathcal{S}(\boldsymbol{\Delta}_{k})$$\triangleq$ $\bigcap_{i=1}^{q}P_{\mathcal{R}_{i}}^{-1}\left(\mathcal{B}_{i}\left(\Delta_{k,i}\right)\right)=\left\{s\in{\mathbb{R}}^{n}\mid\|s^{\mathcal{I}_{i}}\|\leq\Delta_{k,i},\ i=1,2,\ldots,q\right\},$
where $\boldsymbol{\Delta}_{k}=[\Delta_{k,1},\ldots,\Delta_{k,q}]$. As before, $\|s^{\mathcal{I}_{i}}\|$ is shorthand for $\|P_{\mathcal{R}_{i}}(s)\|$. When the objective takes the form $f(x,y,z)=f_{1}(x,z)+f_{2}(y,z)$, $\|s^{\mathcal{I}_{i}}\|$ uses the $L_{2}$-norm, and $\Delta_{k,1}=\Delta_{k,2}\triangleq\overline{\Delta}$, the set $\mathcal{S}(\boldsymbol{\Delta}_{k})$ will reduce to a Steinmetz solid, which is the intersection of two right circular cylinders with equal radii and orthogonal axes, as illustrated in Figure 1.
[…]
Fully described in the text.
Figure 1. Generation of the surface of a Steinmetz solid.
When $\Delta_{k,1}\neq\Delta_{k,2}$, the shape of $\mathcal{S}(\boldsymbol{\Delta}_{k})$ becomes elongated, indicating that the model is more reliable in one direction than another, thus permitting larger steps in the reliable direction. In more general cases, the geometry of $\mathcal{S}(\boldsymbol{\Delta}_{k})$ is highly complex, making the trust-region problem
(6) $\min\limits_{s\in\mathcal{S}(\boldsymbol{\Delta}_{k})}$ $m_{k}(x+s)=\sum_{i=1}^{q}m_{k,i}(x+s),$
rather challenging to solve. It is straightforward to observe that the cylinder set $\mathcal{S}(\boldsymbol{\Delta}_{k})$ contains an inscribed ball with radius $\Delta_{k,\min}\triangleq\min\limits_{i=1,\ldots,q}\Delta_{k,i}$, denoted as $\mathcal{B}(\Delta_{k,\min})$. Thus, a cheap alternative is to solve the problem (6) approximately within $\mathcal{B}(\Delta_{k,\min})$. To address (6), Conn et al. (Conn et al., 1996) proposed a two-stage strategy. The first stage involves minimizing $m_{k}\left(x_{k}-t\nabla m_{k}(x_{k})\right)$ within $\mathcal{B}(\Delta_{k,\min})$ to obtain a Cauchy step $s_{C}$. If further reduction can be achieved by increasing the step length, a subsequent minimization of $m_{k}\left(x_{k}-ts_{C}\right)$ is then attempted within $\mathcal{S}(\boldsymbol{\Delta}_{k})$. In another DFO method PSDFO (Colson and Toint, 2005), all element models share the same trust-region radius, thereby circumventing the difficulty of handling structured trust regions. If derivatives are available, it is reasonable to reduce computational costs by simplifying the trust region geometry—for instance, replacing the original region with an inscribed ellipsoid or employing the $L_{\infty}$-norm in the constraints to transform the trust region into a cube (Shahabuddin, 1996). Nevertheless, to avoid potential degradation in subproblem solution accuracy caused by such simplifications, we prefer to maintain the trust region $\mathcal{S}(\boldsymbol{\Delta}_{k})$ in its original form.
Since $\mathcal{S}(\boldsymbol{\Delta}_{k})$ is the intersection of finitely many convex sets, it remains convex. To this end, we design an approximate projection operator $\widehat{P}_{\mathcal{S}(\boldsymbol{\Delta}_{k})}\colon{\mathbb{R}}^{n}\to\mathcal{S}(\boldsymbol{\Delta}_{k})$ and employ a modified projected gradient method to solve problem (6). The notation $P_{\mathcal{S}(\boldsymbol{\Delta}_{k})}$ is reserved for the mathematically exact projection.
For projections onto convex sets formed by the intersection of multiple convex sets, a popular approach is to use the averaged projection (Lewis et al., 2007). Given arbitrary closed convex sets $F_{1},\ldots,F_{q}$, the averaged projection is defined as
$\overline{P}_{F_{1},\ldots,F_{q}}\triangleq\frac{1}{q}\sum_{i=1}^{q}P_{F_{i}},$
where $P_{F_{i}}$ is the projection operator from $\mathbb{R}^{n}$ to $F_{i}$. For any point $x_{1}$ to be projected, one can iteratively compute $x_{k+1}=\overline{P}_{F_{1},\ldots,F_{q}}(x_{k})$ for $k=1,2,\ldots$ to obtain an approximation of the projection $P_{\mathcal{S}(\boldsymbol{\Delta}_{k})}(x_{1})$. Lewis et al. (Lewis et al., 2007) proved that, under a set of mild conditions on $F_{1},\ldots,F_{q}$, this iteration reduces the mean squared distance $\sum_{i=1}^{q}d(x_{k},F_{i})^{2}/q$ at a linear rate of $1-c/q$, where $c$ is a constant depending on $F_{1},\ldots,F_{q}$. This convergence rate implies that $O\left(\log\varepsilon/\log(1-c/q)\right)$ iterations are sufficient to reduce the mean squared distance below any given tolerance $\varepsilon>0$. Since $\log(1-c/q)<-c/q$ always holds and $\log(1-c/q)$ approaches $-c/q$ when $q$ is large, the required number of iterations scales as $O(q\log(1/\varepsilon))$.
Each iteration of averaged projection involves recomputing all projections $P_{F_{1}}(x_{k}),\ldots,P_{F_{q}}(x_{k})$. When $F_{i}$ is a ball or a cylinder, computing $P_{F_{i}}(x_{k})$ costs $O(n)$. Thus, the overall computational complexity of the averaged projection method is $O(nq^{2}\log(1/\varepsilon))$.
Input: Point $s_{0}\in\mathbb{R}^{n}$, index sets $\{\mathcal{I}_{i}\}_{i=1}^{q}$, trust-region radii $\{\Delta_{i}\}_{i=1}^{q}$
Output: An approximation of $P_{\mathcal{S}(\boldsymbol{\Delta})}(s_{0})$
1
2for $k=0,1,\ldots$ do
3       Compute $v_{i}\leftarrow\left\|s_{0}^{\mathcal{I}_{i}}\right\|_{2}/\Delta_{i}$ for $i=1,\ldots,q$;
4
5      if $\max\limits_{i=1,\ldots,q}v_{k,i}\leq 1$ then
6             Break
7
8      $\mathcal{G}_{k}\leftarrow\left\{i=1,\ldots,q\mid v_{k,i}=\max\limits_{i=1,\ldots,q}v_{k,i}\right\}$;
9
10      $s_{k}\leftarrow\texttt{SHRINK}\left(s_{k},~{}\{\mathcal{I}_{i}\}_{i=1}^{q},~{}\{\Delta_{i}\}_{i=1}^{q},~{}\mathcal{G}_{k}\right)$;
11
Algorithm 2 Steinmetz Projection
To improve the computational efficiency of projection operations, we propose an alternative approximate projection operator tailored for convex sets of the form $\mathcal{S}(\boldsymbol{\Delta})$, referred to as the Steinmetz projection. The procedure is outlined in Algorithm 2, which employs a subroutine SHRINK (see Algorithm 3). For clarity, we define the violation ratio of a point $s\in\mathbb{R}^{n}$ with respect to the $i$-th element as
$v_{i}(s)\triangleq\frac{\left\|s^{\mathcal{I}_{i}}\right\|_{2}}{\Delta_{i}}.$
The core idea of Steinmetz projection stems from an observation: Given an index set $\mathcal{G}$ and a point $s\in\mathbb{R}^{n}$ where each projection $s^{\mathcal{I}_{i}}$ satisfies $\|s^{\mathcal{I}_{i}}\|=\mu\Delta_{i}$ for all $i\in\mathcal{G}$, then scaling $s$ by any real factor $t$ yields $\|(ts)^{\mathcal{I}_{i}}\|=t\mu\Delta_{i}$ for all $i\in\mathcal{G}$. Thus, when projecting $s$ onto $\mathcal{S}(\boldsymbol{\Delta})$, we first collect the indices of all elements with the highest violation ratios into a set $\mathcal{G}$, which by construction satisfies $\|s^{\mathcal{I}_{i}}\|=\mu\Delta_{i}$ for some $\mu\in\mathbb{R}$ and all $i\in\mathcal{G}$. Next, we rescale $s$ in the subspaces corresponding to $\mathcal{G}$ until another element can be added into $\mathcal{G}$, repeating the process iteratively. Here, $\mathcal{G}$ is termed the shrinking set. Once $\mathcal{G}$ encompasses all elements and the current point $s$ satisfies $\|s^{\mathcal{I}_{i}}\|_{2}=\mu\Delta_{i}$ for all $i$, a final scaling by $1/\mu$ ensures the resulting point lies within $\mathcal{S}(\boldsymbol{\Delta})$. As shown in Algorithm 2, each iteration first identifies the shrinking set and then performs the shrinking operation.
Input: Point $s\in\mathbb{R}^{n}$, index sets $\{\mathcal{I}_{i}\}_{i=1}^{q}$, trust-region radii $\{\Delta_{i}\}_{i=1}^{q}$, shrinking set $\mathcal{G}$
Output: Shrunk point $s_{*}$
1
2Select arbitrary $i\in\mathcal{G}$ and compute $u\leftarrow\left\|s^{\mathcal{I}_{i}}\right\|_{2}/\Delta_{i}$;
3
4if $\mathcal{G}=\{1,2,\ldots,q\}$ then
5       $s_{*}\leftarrow s/u$;
6
7else
8       $\mathcal{I}_{\mathcal{G}}\leftarrow\bigcup_{i\in\mathcal{G}}\mathcal{I}_{i}$;
9       for $i=1,\ldots,q$ do
10             if $i\in\mathcal{G}$ then
11                   $t_{i}\leftarrow 1/u$;
12
13            else
14                   $t_{i}\leftarrow\max\left\{\frac{1}{u},\ \frac{\|s^{\mathcal{I}_{i}\backslash\mathcal{I}_{\mathcal{G}}}\|_{2}}{\sqrt{u^{2}\Delta_{i}^{2}-\|s^{\mathcal{I}_{i}\cap\mathcal{I}_{\mathcal{G}}}\|_{2}^{2}}}\right\}$;
15
16
17      $t_{*}\leftarrow\max_{i=1,\ldots,q}t_{i}$;
18       $s_{*}\leftarrow s+\left(t_{*}-1\right)s^{\mathcal{I}_{\mathcal{G}}}$;
19
Algorithm 3 Subroutine SHRINK
The SHRINK subroutine (Algorithm 3) implements the shrinking operation. Assuming without loss of generality that the given point $s$ lies outside $\mathcal{S}(\boldsymbol{\Delta})$, and given a shrinking set $\mathcal{G}=\{1,2,\ldots,\hat{q}\}\ (\hat{q}<q)$ satisfying
(7) $v_{1}(s)=\cdots=v_{\hat{q}}(s)>v_{i}(s),\quad\forall i\in\{\hat{q}+1,\ldots,q\},$
the subroutine performs component-wise contraction in the subspace $\mathcal{R}_{\mathcal{G}}\triangleq\operatorname{span}\left(\cup_{i\in\mathcal{G}}\mathcal{R}_{i}\right)$. Specifically, it scales the component $s^{\mathcal{I}_{\mathcal{G}}}$ by a factor $t_{*}\in\mathbb{R}$, generating a new point $s_{*}\triangleq s+(t_{*}-1)s^{\mathcal{I}_{\mathcal{G}}}$ that satisfies either:
(8) $1=v_{1}(s_{*})=\cdots=v_{\hat{q}}(s_{*})\geq v_{i}(s_{*}),\quad\forall i\in\{\hat{q}+1,\ldots,q\},$
or introduces an elemental index $\hat{p}\in\{\hat{q}+1,\ldots,q\}$ such that
(9) $v_{1}(s_{*})=\cdots=v_{\hat{q}}(s_{*})=v_{\hat{p}}(s_{*})\geq v_{i}(s_{*}),\quad\forall i\in\{\hat{q}+1,\ldots,q\}\setminus\{\hat{p}\},$
where $\mathcal{I}_{\mathcal{G}}\triangleq\bigcup_{i\in\mathcal{G}}\mathcal{I}_{i}.$
This mechanism enables the outer loop (Algorithm 2) to terminate the projection or subsequently expand the shrinking set to $\mathcal{G}_{\text{new}}=\{1,2,\ldots,\hat{q}\}\cup\{\hat{p}\}$. The elemental scaling ratio $t_{i}$ is the maximum of two possible choices, as computed in Steps 3 and 3 of Algorithm 3, yielding the candidate point $s_{i}\triangleq s+(t_{i}-1)s^{\mathcal{I}_{\mathcal{G}}}$ satisfying either $1=v_{1}(s_{i})=\cdots=v_{\hat{q}}(s_{i})>v_{i}(s_{i})$ or $v_{1}(s_{i})=\cdots=v_{\hat{q}}(s_{i})=v_{i}(s_{i})$. The final contraction ratio $t_{*}$ is then selected as the maximum value of $t_{1},\ldots,t_{q}$ ensuring satisfaction of (8) or (9). Note that the selection of trial index $i$ in Step 3 is arbitrary, as the ratio $\left\|s^{\mathcal{I}_{i}}\right\|_{2}/\Delta_{i}$ remains constant across all $i\in\mathcal{G}$ due to the defining property of the shrinking set $\mathcal{G}$. Theorem 2.1 formally establishes the correctness of the SHRINK subroutine.
Theorem 2.1.
Given a point $s\notin\mathcal{S}(\boldsymbol{\Delta})$, element index sets $\{\mathcal{I}_{i}\}_{i=1}^{q}$, trust-region radii $\{\Delta_{i}\}_{i=1}^{q}$, and a shrinking set $\mathcal{G}$ satisfying condition (7), there exists $\hat{p}\in\{\hat{q}+1,\ldots,q\}$ such that the point $s_{*}$ returned by Algorithm 3 will satisfy either (8) or (9).
Proof.
Let $\hat{p}=\operatorname{argmax}_{i\notin\mathcal{G}}t_{i}$, where $t_{i}$ is determined by Steps 3 and 3 of Algorithm 3. Let $u=\left\|s^{\mathcal{I}_{i}}\right\|_{2}/\Delta_{i}$ for an arbitrary $i\in\mathcal{G}$, noting that this value is independent of the choice of $i$. Define $t_{*}=\max_{i=1,\ldots,q}t_{i}$ and $s_{*}=s+(t_{*}-1)s^{\mathcal{I}_{\mathcal{G}}}$. We consider the size of $u^{2}\Delta_{p}^{2}$ in two cases.
First, if $u^{2}\Delta_{\hat{p}}^{2}>u^{2}\left\|s^{\mathcal{I}_{\hat{p}}\backslash\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}+\left\|s^{\mathcal{I}_{\hat{p}}\cap\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}$ holds, then we have
$\frac{1}{u}>\frac{\|s^{\mathcal{I}_{i}\backslash\mathcal{I}_{\mathcal{G}}}\|_{2}}{\sqrt{u^{2}\Delta_{i}^{2}-\|s^{\mathcal{I}_{i}\cap\mathcal{I}_{\mathcal{G}}}\|_{2}^{2}}},$
and thus $t_{\hat{p}}=1/u$. On one hand, $t_{i}=1/u~{}(\forall i\in\mathcal{G})$, and on the other hand, $t_{i}\leq t_{\hat{q}}=1/u~{}(\forall i\notin\mathcal{G})$. Thus, $t_{*}=1/u$. For any $i\in\mathcal{G}$, since $\mathcal{I}_{i}\subset\mathcal{I}_{\mathcal{G}}$, we have
(10) $s_{*}^{\mathcal{I}_{i}}=P_{\mathcal{R}_{i}}s_{*}=t_{*}s^{\mathcal{I}_{i}},\quad\forall i\in\mathcal{G}.$
Furthermore, by the definition of $u$, it follows that
$s_{*}^{\mathcal{I}_{i}}=\frac{1}{u}s^{\mathcal{I}_{i}}\Longrightarrow\left\|s_{*}^{\mathcal{I}_{i}}\right\|_{2}=\frac{1}{u}\left\|s^{\mathcal{I}_{i}}\right\|_{2}=\Delta_{i}\Longrightarrow v_{i}(s_{*})=1,\quad$ $\forall i\in\mathcal{G}.$
Meanwhile, for any $i\notin\mathcal{G}$, since
(11) $s_{*}^{\mathcal{I}_{i}}=s^{\mathcal{I}_{i}\backslash\mathcal{I}_{\mathcal{G}}}+t_{*}s^{\mathcal{I}_{i}\cap\mathcal{I}_{\mathcal{G}}},\quad\forall i\notin\mathcal{G},$
and $s^{\mathcal{I}_{i}\backslash\mathcal{I}_{\mathcal{G}}}\perp s^{\mathcal{I}_{i}\cap\mathcal{I}_{\mathcal{G}}}$, we have
$\left\|s_{*}^{\mathcal{I}_{i}}\right\|_{2}^{2}=\left\|s^{\mathcal{I}_{i}\backslash\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}+\frac{1}{u^{2}}\left\|s^{\mathcal{I}_{i}\cap\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}\leq\Delta_{i}^{2}\Longrightarrow v_{i}(s_{*})\leq 1,\quad\forall i\notin\mathcal{G}.$
Thus, in this case, inequality (8) holds.
In the second case, when $u^{2}\Delta_{\hat{p}}^{2}\leq u^{2}\left\|s^{\mathcal{I}_{\hat{p}}\backslash\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}+\left\|s^{\mathcal{I}_{\hat{p}}\cap\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}$ holds, we have $t_{\hat{p}}\geq 1/u$, and thus $t_{*}=t_{\hat{p}}$. For any $i\in\mathcal{G}$, since equation (10) still holds, it follows that $v_{i}(s_{*})=t_{*}v_{i}(s)=t_{\hat{p}}u$. Therefore, all $v_{i}(s_{*})$ ($\forall i\in\mathcal{G}$) remain the same. From equation (11), we derive
$\left\|s_{*}^{\mathcal{I}_{\hat{p}}}\right\|_{2}^{2}=\left\|s^{\mathcal{I}_{\hat{p}}\backslash\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}+t_{\hat{p}}^{2}\left\|s^{\mathcal{I}_{\hat{p}}\cap\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}=\left\|s^{\mathcal{I}_{\hat{p}}\backslash\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}+\frac{\left\|s^{\mathcal{I}_{\hat{p}}\backslash\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}\left\|s^{\mathcal{I}_{\hat{p}}\cap\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}}{u^{2}\Delta_{\hat{p}}^{2}-\left\|s^{\mathcal{I}_{\hat{p}}\cap\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}}=\frac{u^{2}\Delta_{\hat{p}}^{2}\left\|s^{\mathcal{I}_{\hat{p}}\backslash\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}}{u^{2}\Delta_{\hat{p}}^{2}-\left\|s^{\mathcal{I}_{\hat{p}}\cap\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}}=u^{2}\Delta_{\hat{p}}^{2}t_{\hat{p}}^{2}.$
Thus, we have $v_{\hat{p}}(s_{*})=t_{\hat{p}}u$. For any $i\notin\mathcal{G}$, we claim that $v_{i}(s_{*})\leq t_{\hat{p}}u$. Otherwise, the inequality $\left\|s_{*}^{\mathcal{I}_{i}}\right\|_{2}>t_{\hat{p}}u\Delta_{i}$ would hold, leading to
$\left\|s^{\mathcal{I}_{i}\backslash\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}+t_{\hat{p}}^{2}\left\|s^{\mathcal{I}_{i}\cap\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}>t_{\hat{p}}^{2}u^{2}\Delta_{i}^{2}\Longrightarrow\frac{\left\|s^{\mathcal{I}_{i}\backslash\mathcal{I}_{\mathcal{G}}}\right\|_{2}}{\sqrt{u^{2}\Delta_{i}^{2}-\left\|s^{\mathcal{I}_{i}\cap\mathcal{I}_{\mathcal{G}}}\right\|_{2}^{2}}}>t_{\hat{p}},$
which implies $t_{i}>t_{\hat{p}}$. This contradicts the assumption $\hat{p}=\operatorname{argmax}_{i\notin\mathcal{G}}t_{i}$. Therefore, we have
$t_{\hat{p}}u=v_{j}(s_{*})=v_{\hat{p}}(s_{*})\geq v_{i}(s_{*}),\quad\forall~{}i\in\mathcal{G}^{c}\backslash\{\hat{p}\},~{}j\in\mathcal{G}.$
Hence, inequality (9) holds. ∎
Note that at the initialization of Steinmetz projection, we may exclude the index sets $\mathcal{I}_{j}$ of elements satisfying $v_{j}(s_{0})\leq 1$ from $\{\mathcal{I}_{i}\}_{i=1}^{q}$, since they never contribute $t_{j}$ values exceeding $1/u$ and thus cannot affect the value of $t_{*}$. Furthermore, Steps 2 and 2 of Algorithm 2 need not be explicitly recomputed in each iteration, as the new shrinking set can always be obtained simply using intermediate results from the subroutine SHRINK. If |$\left|\left\{i\mid v_{j}(s_{0})\leq 1\right\}\right|=\overline{q}$, then each run of Algorithm 3 has a cost of $O\left(n\left(q-\overline{q}-|\mathcal{G}|\right)\right)$. Since each run either terminates the projection or adds at least one new member to the shrinking set, the algorithm iterates at most $q-\overline{q}$ times. Therefore, the computational complexity of Algorithm 2 is
(12) $O\left(\sum_{i=1}^{q-\overline{q}}n(q-\overline{q}-i)\right)=O\left(\frac{1}{2}n(q-\overline{q})^{2}\right).$
Meanwhile, the averaged projection method with stopping tolerance $\varepsilon$ has a cost of $O(n(q-\overline{q})^{2}\log\varepsilon)$. By contrast, the Steinmetz projection does not require setting termination precision, resulting in a lower constant factor. Figure 2 illustrates the iterates generated by applying both averaged projection and Steinmetz projection to the initial point $s_{0}$, where the labeled $\mathcal{S}(\boldsymbol{\Delta})$, $P_{\mathcal{R}_{1}}^{-1}\left(\mathcal{B}(\Delta_{1})\right)$, and $P_{\mathcal{R}_{2}}^{-1}\left(\mathcal{B}(\Delta_{2})\right)$ represent their cross-sections in the current region.
[…]
[The averaged projection method may take much more steps to converge than the Steinmetz projection method.] Iterates generated by the two projection methods, of which $s_{0},s_{1},s_{2}$ are the iterates generated by the Steinmetz projection, $s_{0},s^{\prime}_{1},\ldots,s^{\prime}_{\infty}$ are the iterates generated by the averaged projection, and $s_{2},s^{\prime}_{\infty}$ are the finishing points. The averaged projection method may take much more steps to reach the finishing point than the Steinmetz projection method.
Figure 2. Iterates generated by applying the averaged projection and the Steinmetz projection to the initial point $s_{0}$, where the marked $\mathcal{S}(\boldsymbol{\Delta})$, $P_{\mathcal{R}_{1}}^{-1}(\mathcal{B}(\Delta_{1}))$, and $P_{\mathcal{R}_{2}}^{-1}(\mathcal{B}(\Delta_{2}))$ represent their cross-sections in the current region.
It should be noted that neither the Steinmetz projection nor the averaged projection method constitutes a projection in the mathematical sense (i.e., finding the closest point under a given norm), and it is difficult to assert which method consistently yields better optimality for the projected points. Based on our numerical tests, we found that first applying the averaged projection for $k_{\text{avg}}=1\sim 5$ iterations, followed by the Steinmetz projection, sometimes produces a point closer to the initial point than using either method alone. For any point $s\in{\mathbb{R}}^{n}$ to be projected, we denote the resulting projected point obtained through this combined approach as $\widehat{P}_{\mathcal{S}(\boldsymbol{\Delta})}\left(s;~{}k_{\text{avg}}\right)$.
To solve the trust-region subproblem (6), we employ a modified projected gradient method (Algorithm 4), which closely follows the truncated conjugate gradient method. The key difference arises when the iterate $s_{k}+\alpha_{k}p_{k}$ falls outside $\mathcal{S}(\boldsymbol{\Delta})$. Instead of stopping, the algorithm performs an exact line search along the direction $\widehat{P}_{\mathcal{S}(\boldsymbol{\Delta})}(s_{k}+\alpha_{k}p_{k};~{}k_{\text{avg}})-s_{k}$ starting from $s_{k}$ to determine $s_{k+1}$. Since $\widehat{P}_{\mathcal{S}(\boldsymbol{\Delta})}(\cdot;~{}k_{\text{avg}})$ is not an exact projection, monotonic decrease in the function values cannot be guaranteed, making this additional line search necessary. Furthermore, the algorithm alternates between $k_{\text{avg}}^{(1)}=0$ and $k_{\text{avg}}^{(2)}=4$ during iterations to mitigate the limitations of using a single projection operator. The truncation radius $\widehat{\Delta}$ in Algorithm 4 is chosen sufficiently large to ensure that the trust region is entirely contained within the ball $\mathcal{B}(\widehat{\Delta})$. In addition, a comparison of UPOQA’s performance with versus without structured trust regions is provided in Appendix B, where the non-structured version employs a spherical trust region, and the resulting subproblem is solved via the truncated conjugate gradient method with a boundary improvement step.
Input: Overall model $m$, center point $x$, truncation radius $\widehat{\Delta}>0$, projection operator $\widehat{P}_{\mathcal{S}(\boldsymbol{\Delta})}$, numbers of start-up averaged projection iterations $k_{\text{avg}}^{(1)},~{}k_{\text{avg}}^{(2)}\geq 0$
Output: An approximate solution to (6)
1 Set the iterate $s_{0}\leftarrow 0$;
2 Set the search direction $p_{0}\leftarrow-\nabla m(x+s_{0})$;
3 for $k=0,1,\ldots~{}\mathbf{until}~{}\nabla m(x+s_{k})^{\top}p_{k}\geq 0$  do
4
5      Compute $\alpha_{k}\leftarrow\mathop{\text{argmin}}\left\{m(x+s_{k}+\alpha p_{k})\mid 0\leq\alpha\leq\widehat{\Delta}/\|p_{k}\|_{2}\right\}$;
6
7      if $s_{k}+\alpha^{k}p_{k}$ falls outside $\mathcal{S}(\boldsymbol{\Delta})$ then
8             $\hat{s}_{k}\leftarrow\widehat{P}_{\mathcal{S}(\boldsymbol{\Delta})}\left(s_{k}+\alpha_{k}p_{k};~{}k_{\text{avg}}^{(k~{}\mathrm{mod}~{}2)}\right)$;
9             $\hat{p}_{k}\leftarrow\hat{s}_{k}-s_{k}$;
10             $\hat{\alpha}_{k}\leftarrow\mathop{\text{argmin}}\left\{m(x+s_{k}+\alpha\hat{p}_{k})\mid 0\leq\alpha\leq 1\right\}$;
11             $s_{k+1}\leftarrow s_{k}+\hat{\alpha}_{k}\hat{p}_{k}$;
12             $p_{k+1}\leftarrow-\nabla m(x+s_{k+1})$;
13
14      else
15             $s_{k+1}\leftarrow s_{k}+\alpha_{k}p_{k}$;
16             $\beta_{k}\leftarrow\left\|\nabla m(x+s_{k+1})\right\|^{2}/\left\|\nabla m(x+s_{k})\right\|^{2}$;
17             $p_{k+1}\leftarrow-\nabla m(x+s_{k+1})+\beta_{k}p_{k}$;
18
19
20
Algorithm 4 Modified Projected Gradient Method for the Trust-Region Subproblem (6)
2.2. Managment of Trust Region Radii
In Step 1 of Algorithm 1, we employ the combined separation criterion proposed by Shahabuddin (Shahabuddin, 1996, Section 2.4) to adjust the trust region radii $\{\Delta_{k,i}\}_{i=1}^{q}$. Here we briefly introduce this criterion.
In trust region methods, it is common practice to adjust the trust region based on the value of the reduction ratio $r_{k}=\delta_{k}f/\delta_{k}m_{k}$, where
$\delta_{k}f=f(x_{k})-f(\hat{x}_{k}),\quad\delta_{k}m_{k}=m_{k}(x_{k})-m_{k}(\hat{x}_{k}).$
However, when solving partially separable problems, it is infeasible to independently adjust each $\Delta_{k,i}$ solely based on $r_{k,i}$. For an individual objective, we always expect the surrogate model and the objective function to decrease. However, for coexisting element functions, the potential antagonistic relationship between them may necessitate that we sometimes tolerate a slight deterioration in the function values of some elements. This suggests that the algorithm should not simply encourage larger $r_{k,i}$ values. Moreover, the cancellation issue further complicates the problem. Conn et al. (Conn et al., 1996) first pointed out that a situation may arise where $\delta_{k}m_{k,1},\ldots,\delta_{k}m_{k,q}$ have large absolute magnitudes, but cancellation occurs, resulting in a very small absolute value for the aggregated $\delta_{k}m_{k}$. In such cases, rounding errors in $\delta_{k}m_{k,1},\ldots,\delta_{k}m_{k,q}$ can easily lead to inaccurate computation of $\delta_{k}m_{k}$ and to misjudgment of the reliability of the overall model $m_{k}$.
For above considerations, Shahabuddin (Shahabuddin, 1996, Section 2.4) proposed a scoring-based strategy. The scores consist of a global score $\tau_{k}$ and element-wise local scores $\{\hat{\tau}_{k,i}\}_{i=1}^{q}$, both taking integer values of 0, 1, or 2. By summing the two parts, the total score (ranging from 0 to 4) determines how each element’s trust region radius $\Delta_{k,i}$ should be adjusted. The global score $\tau_{k}$ depends solely on the magnitude of $r_{k}$—the larger $r_{k}$, the higher $\tau_{k}$. The element-wise score $\hat{\tau}_{k,i}$ is determined by the position of the point $(\delta_{k}m_{k,i},\delta_{k}f_{i})$. Figure 3 illustrates the acceptable region for assigning $\{\hat{\tau}_{k,i}\}_{i=1}^{q}$, bounded by two rays with different slopes emanating from the origin and a line with slope 1 that does not pass through the origin. The stricter the acceptable region in which $(\delta_{k}m_{k,i},\delta_{k}f_{i})$ falls, the higher $\hat{\tau}_{k,i}$. The detailed scoring strategy is described in Algorithm 5. After scoring, we set $\tau_{k,i}=\tau_{k}+\hat{\tau}_{k,i}$. The new trust region radii are then updated as follows:
$\Delta_{k+1,i}\begin{cases}\in[1,~{}\theta_{4}]\Delta_{k,i},&\text{if }~{}\tau_{k,i}=4,\\ \in[1,~{}\theta_{3}]\Delta_{k,i},&\text{if }~{}\tau_{k,i}=3,\\ \in[\theta_{2},~{}1]\Delta_{k,i},&\text{if }~{}\tau_{k,i}=2,\\ =\theta_{2}\Delta_{k,i},&\text{if }~{}\tau_{k,i}=1,\\ =\theta_{1}\Delta_{k,i},&\text{if }~{}\tau_{k,i}=0,\end{cases}$
In our implementation, the parameters are set as $\theta_{1}=0.5$, $\theta_{2}=1/\sqrt{2}$, $\theta_{3}=\sqrt{2}$, and $\theta_{4}=2$. Additionally, Step 5 in Algorithm 5 ensures that when the overall model reduction is unsatisfactory ($r_{k}<\mu_{1}$), at least one element’s trust region radius undergoes substantial decrease to prevent the algorithm from stagnating at a fixed radius. Although such cases are rare, they have indeed been observed in our numerical experiments. This step is the only modification we made to Shahabuddin’s combined separation criterion (Shahabuddin, 1996, Section 2.4).
Figure 3. The combined separation criterion proposed by Shahabuddin (Shahabuddin, 1996, Section 2.4) determines whether to expand the elemental trust region radius $\Delta_{k,i}$ based on the slope $r_{k,i}$ of the point $(\delta_{k}m_{k,i},~{}\delta_{k}f_{i})$ relative to the origin and its distance to the line $r_{k,i}=1$. The colored region in the figure represents the acceptable region for expansion operations.
[…]
Fully described in the text.
Input: Elemental model reductions $\{\delta_{k}m_{k,i}\}_{i=1}^{q}$, elemental reduction ratios $\{r_{k,i}\}_{i=1}^{q}$, trust region radii $\{\Delta_{k,i}\}_{i=1}^{q}$, trust region resolution $\rho_{k}$, overall reduction ratio $r_{k}$, constants $10<\mu_{1}<\mu_{2}<$ 1
Output: Total scores for each element $\{\tau_{k,i}\}_{i=1}^{q}$
1
[…]
3
4For $i=1,2$, set $\eta_{i}=-(1-\mu_{i})\zeta_{k}$, $\alpha_{i}\leftarrow\frac{(\mu_{i}+\eta_{i})(1+\zeta_{k})-2\zeta_{k}}{1-\zeta_{k}}$;
5
6if $r_{k}\geq\mu_{2}$ then  $\tau_{k}\leftarrow 2$ ;
7else if $r_{k}\geq\mu_{1}$ then  $\tau_{k}\leftarrow 1$ ;
8else  $\tau_{k}\leftarrow 0$ ;
9
10for $i=1,2,\ldots,q$ do
11       if $\delta_{k}m_{k,i}\geq 0$ then
12             if $r_{k,i}\geq\alpha_{2}~{}\mathbf{or}~{}\delta_{k}f_{i}\geq\delta_{k}m_{k,i}-\eta_{2}(\delta_{k}m_{k})/q$  then
13                   $\hat{\tau}_{k,i}\leftarrow 2$;
14
15            else if $r_{k,i}\geq\alpha_{1}~{}\mathbf{or}~{}\delta_{k}f_{i}\geq\delta_{k}m_{k,i}-\eta_{1}(\delta_{k}m_{k})/q$ then
16                   $\hat{\tau}_{k,i}\leftarrow 1$;
17
18             else  $\hat{\tau}_{k,i}\leftarrow 0$ ;
19
20      else
21             if $r_{k,i}\leq 2-\alpha_{2}~{}\mathbf{or}~{}\delta_{k}f_{i}\geq\delta_{k}m_{k,i}-\eta_{2}(\delta_{k}m_{k})/q$ then
22                   $\hat{\tau}_{k,i}\leftarrow 2$;
23
24            else if $r_{k,i}\leq 2-\alpha_{1}~{}\mathbf{or}~{}\delta_{k}f_{i}\geq\delta_{k}m_{k,i}-\eta_{1}(\delta_{k}m_{k})/q$ then
25                   $\hat{\tau}_{k,i}\leftarrow 1$;
26
27            else  $\hat{\tau}_{k,i}\leftarrow 0$ ;
28
29      $\tau_{k,i}\leftarrow\hat{\tau}_{k,i}+\tau_{k}$;
30
31if $\tau_{k}=0$ then
32       Verify all $\tau_{k,i}$ to ensure at least one element satisfies $\tau_{k,i}\leq 1$ and $\Delta_{k,i}>\rho_{k}$. Otherwise, set $\tau_{k,i}$ of the lowest-scoring element with $\Delta_{k,i}>\rho_{k}$ to 00;
33
Algorithm 5 Scoring Strategy for Trust Region Radius Adjustment
2.3. Selective Updates of Element Models
After obtaining the solution $s_{k}$ to the trust-region subproblem (6), it is time to update element models and interpolation sets. For the model $m_{k,i}$, the new interpolation point generated by the solution $s_{k}$ is $\hat{x}_{k}^{\mathcal{I}_{i}}=(x_{k}+s_{k})^{\mathcal{I}_{i}}$. However, blindly adding $\hat{x}_{k}^{\mathcal{I}_{i}}$ to the set $\mathcal{Y}_{k,i}$ may deteriorate the poisedness of $\mathcal{Y}_{k,i}$. For example, when the objective function is $f(x,y)=f_{1}(x)+f_{2}(y)$ and the starting point is $(x_{0},y_{0})$, if $x_{0}$ is already very close to a local minimizer of $f_{1}$, the computed trust-region step will have vastly different scales in the subspaces corresponding to $x$ and $y$. This can cause a drastic contraction in the geometry of the interpolation set $\mathcal{Y}_{1}$, potentially leading to numerical instability. If the algorithm detects such a risk, it skips such updates, as implemented in Step 1 of Algorithm 1.
In Powell’s DFO methods that based on quadratic models, the denominator $\sigma$ in the update formula for the interpolation system is a crucial parameter (Powell, 2004b). Its magnitude largely determines the impact of the update on the poisedness of the interpolation system, and a larger value is preferred. Therefore, we decide whether to accept the new interpolation point $\hat{x}_{k}^{\mathcal{I}_{i}}$ into $\mathcal{Y}_{k,i}$ by checking the value $\sigma_{k,i}$ during the update of $m_{k,i}$. We first compute $\sigma_{k,i}$ and multiply it by a factor $\gamma_{k,i}\triangleq\min\left\{\left\|s_{k}^{\mathcal{I}_{i}}\right\|_{2}/\rho_{k},~{}1\right\}$, where $s_{k}$ is the solution to the trust-region subproblem (6) solved in Step 8 of Algorithm 1. The factor $\gamma_{k,i}$ penalizes situations with small step updates, as they generally contribute limited improvements to the model’s accuracy, even if they don’t significantly compromise the system’s poisedness. Due to rounding errors, the value of $\sigma_{k,i}$ may become negative. To mitigate this, we check the intermediate variables used in computing $\sigma_{k,i}$ to detect such tendencies, and apply an additional penalty factor $w_{k,i}$ to the value of $\sigma_{k,i}$. This penalty factor is determined heuristically - in short, we reduce positive $\sigma_{k,i}$ values while enlarging negative ones. The resulting value $w_{k,i}\gamma_{k,i}\sigma_{k,i}$ is then compared with a fixed threshold $\xi$. The new interpolation point $\hat{x}_{k}^{\mathcal{I}_{i}}$ is accepted only if $w_{k,i}\gamma_{k,i}\sigma_{k,i}$ exceeds the threshold. In UPOQA, the default threshold is set to $\xi=10^{-5}$.
In most cases, updates with negative $\sigma_{k,i}$ are automatically rejected. However, an exception occurs in extreme situations where all candidate updates yield negative $\sigma_{k,i}$ values, and the algorithm will accept the update whose $w_{k,i}\gamma_{k,i}\sigma_{k,i}$ is closest to zero. Only then does the additional penalty on negative $\sigma_{k,i}$ values matter.
2.4. Starting Point Search
As mentioned earlier, based on the element values evaluated at a certain number of interpolation points selected from the grid
$\left\{x=\sum_{i=1}^{n}c_{i}e_{i}\mid c_{1},\ldots,c_{n}\in\mathbb{Z}\right\},$
one can easily compute significantly more overall function values. This idea has been adopted by Price and Toint (Price and Toint, 2006). Although UPOQA cannot directly utilize this search pattern to accelerate the algorithm, it employs similar strategy to select a better starting point for the algorithm.
UPOQA initializes interpolation sets in the same way as COBYQA (Ragonneau, 2023). Suppose $\mathcal{Y}_{0}=\{y_{0}^{1},\ldots,y_{0}^{N}\}$ is the $n$-dimensional interpolation set to be initialized with a capacity of $N=2n+1$, the initial trust-region radius is $\Delta_{0}$, and the user-specified starting point is $x_{\text{start}}$. Then, the points $y_{0}^{1},\ldots,y_{0}^{N}$ are initialized as
$y_{0}^{i}\triangleq\begin{cases}x_{\text{start}},&\text{if }~{}i=1,\\ x_{\text{start}}+\Delta_{0}e_{i-1},&\text{if }~{}2\leq i\leq n+1,\\ x_{\text{start}}-\Delta_{0}e_{i-n-1},&\text{if }~{}n+2\leq i\leq 2n+1.\end{cases}$
Similarly, for each elemental interpolation set $\mathcal{Y}_{0,i}$, UPOQA initializes it with $x_{\text{start}}^{\mathcal{I}_{i}}$ as the center and $\Delta_{0}$ as the radius. Based on the resulting sets $\{\mathcal{Y}_{0,i}\}_{i=1}^{q}$, UPOQA then explores a certain number of solutions of a constraint satisfaction problem (CSP) using the minimum remaining values heuristic. In this CSP, the state is a complete point $x\in{\mathbb{R}}^{n}$, and the constraints require that for any $i=1,\ldots,q$, $P_{\mathcal{R}_{i}}(x)\in\mathcal{Y}_{0,i}$ must hold. Since the starting point search mechanism has limited impact on improving algorithmic performance, we restrict the search to at most 5000 solutions and select the point with the smallest objective function value as the actual starting point $x_{0}$ for UPOQA.
2.5. Optional Features
Besides the procedure demonstrated in Algorithm 1, UPOQA also provides two additional functionalities to enhance its flexibility and applicability: restart mechanism and hybrid black-white-box optimization.
2.5.1. Restart Mechanism
In model-based derivative-free methods, the trust region radius $\Delta$ or trust region resolution $\rho$ decreasing to a minimal value often indicates algorithm convergence. However, when the objective function contains noise, the noise will dominate the function value fluctuations once $\Delta$ becomes sufficiently small. Consequently, the interpolation model will no longer approximate the objective but model the noise instead. This leads to ineffective descent steps and may even trigger numerical instabilities causing algorithm failure.
Restart mechanisms have been widely adopted in other areas of numerical optimization, such as the conjugate gradient method and GMRES. Cartis et al.’s Py-BOBYQA (Cartis et al., 2019) introduced this mechanism into model-based DFO methods, with experiments demonstrating its robustness for noise. Inspired by this, UPOQA incorporates the soft restart mechanism similar to that in Py-BOBYQA. This mechanism is only activated upon user request. When numerical singularity or trust-region resolution depletion occurs, the algorithm resets $\Delta$ and $\rho$ to larger values, selectively replaces some existing interpolation points with a few randomly generated ones, and relocates the iterate to a distant position before resuming execution.
2.5.2. Hybrid Black-White-Box Optimization
Real-world partially separable problems may exhibit more complex structures than (1). To maximize the algorithm’s applicability, we adapt the implementation of UPOQA to handle problems of the form
(13) $\min_{x\in{\mathbb{R}}^{n}}\ f(x)=f_{0}(x)+\sum_{i=1}^{q}w_{i}h_{i}\left(f_{i}(x)\right),$
where $f_{1},\ldots,f_{q}$ remain black-box functions as before, while $f_{0}\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ is a white-box function whose gradient and Hessian can be calculated directly. Each $h_{i}\colon{\mathbb{R}}\to{\mathbb{R}}$ is a transformation applied to the element value $f_{i}(x)$, also treated as white-box, and $w_{i}\in{\mathbb{R}}$ denotes the weight for each element. The transition from (1) to (13) requires minimal algorithmic modifications, as the added white-box components can be directly incorporated into the trust-region subproblem construction, yet it significantly enhances flexibility. Moreover, UPOQA allows users to dynamically update all $w_{1},\ldots,w_{q}$ and $h_{1},\ldots,h_{q}$ via a callback function during execution. This enables real-time trade-offs among multiple elements in multi-objective optimization, as well as effortless adjustment of penalty coefficients in constrained optimization without manual restarting.
An additional advantage of this adaptation stems from UPOQA’s use of quadratic models for each black-box $f_{i}$. If $f_{i}$ inherently and locally resembles a quadratic function, maintaining an accurate surrogate model becomes computationally inexpensive, reducing unnecessary trial-and-error during optimization. Hence, knowledgeable users may preemptively extract known non-quadratic composite parts from $f_{i}$ and pass them as $h_{i}$ to the algorithm. This strategy lowers the cost of maintaining interpolation models and ultimately decreases the number of function evaluations required for convergence.
3. Numerical Experiments
We briefly describe our testing methodology, the preparation of test problems, and the results of numerical experiments. Our primary evaluation metrics are the performance profile and data profile proposed by Moré and Wild (Moré and Wild, 2009), supplemented with an additional speed-up profile to assess the acceleration effect of UPOQA when exploiting partially separable structures. The UPOQA algorithm has been implemented as a Python library named upoqa. We conducted tests on benchmark problems extracted from CUTEst (Gould et al., 2015) using the S2MPJ (Gratton and Toint, 2024) tool, along with a set of quantum variational problems. All tests were conducted on a server equipped with a 32-core Hygon C86 7285 CPU and 504 GB of RAM.
3.1. Testing Methodology
Let $\mathcal{S}$ denote the set of all solvers participating in the test, and $\mathcal{P}$ denote the set of test problems used. Each test problem $p\in\mathcal{P}$ has a fixed starting point $x_{0,p}$ and an objective function $f_{p}$. Given a tolerance $\varepsilon\in(0,1)$, a solver $s\in\mathcal{S}$ is said to converge on problem $p$ with tolerance $\varepsilon$ if and only if its iterate $x$ satisfies
(14) $f_{p}(x)\leq f_{p}^{*}+\varepsilon\left[f_{p}(x_{0,p})-f_{p}^{*}\right],$
where $f_{p}^{*}$ is the lowest function value achieved by any solver in $\mathcal{S}$ on problem $p$. Let $t_{p,s}$ denote the number of function evaluations required for solver $s$ to converge on problem $p$. If convergence is not achieved, we set $t_{p,s}=\infty$.
The performance profile (Moré and Wild, 2009) is a concept introduced to evaluate and compare different optimization algorithms. This metric assesses which solver has an advantage in computational cost for given test problems. First, the performance ratio is defined as
$r_{p,s}\triangleq\frac{t_{p,s}}{\min_{u\in\mathcal{S}}\{t_{p,u}\}},$
which can be viewed as a relative cost incurred by solver $s$ to solve problem $p$. Then, the performance profile is defined as
$\rho_{s}(\alpha)\triangleq\frac{1}{\operatorname{card}(\mathcal{P})}\operatorname{card}\left(\left\{p\in\mathcal{P}:r_{p,s}\leq\alpha\right\}\right),\quad\text{for}~{}\alpha\geq 1,$
where $\operatorname{card}(\cdot)$ denotes the cardinality of a set. The metric $\rho_{s}(\alpha)$ can be interpreted as the proportion of problems in $\mathcal{P}$ that solver $s$ can solve with a performance ratio not exceeding $\alpha$. Specifically, $\rho_{s}(1)$ represents the proportion of problems where solver $s$ converges faster than all other solvers, while $\rho_{s}(+\infty)$ indicates the proportion of problems that $s$ can solve regardless of cost. Given two solvers $s_{1}$ and $s_{2}$, $\rho_{s_{1}}(\alpha)>\rho_{s_{2}}(\alpha)$ implies that under the constraints $r_{p,s_{1}}\leq\alpha$ and $r_{p,s_{2}}\leq\alpha$, $s_{1}$ can successfully solve more problems than $s_{2}$. The value of $\rho_{s}(\alpha)$ for an individual solver is highly dependent on the competitors present in $\mathcal{S}$.
In contrast, the data profile (Moré and Wild, 2009) focuses more on the absolute capability of a single solver and is defined as
$d_{s}(\alpha)\triangleq\frac{1}{\operatorname{card}(\mathcal{P})}\operatorname{card}\left(\left\{p\in\mathcal{P}:\frac{t_{p,s}}{n_{p}+1}\leq\alpha\right\}\right),\quad\text{for}~{}\alpha\geq 0,$
where $n_{p}$ denotes the dimension of problem $p$. The metric $d_{s}(\alpha)$ can be interpreted as the proportion of problems in $\mathcal{P}$ that solver $s$ can solve using no more than $\alpha(n_{p}+1)$ function evaluations. Specifically, $d_{s}(0)=0$, while $d_{s}(+\infty)$ represents the proportion of problems that $s$ can solve regardless of computational cost. Given two solvers $s_{1}$ and $s_{2}$, $d_{s_{1}}(\alpha)>d_{s_{2}}(\alpha)$ implies that, for each problem $p$, when using no more than $\alpha(n_{p}+1)$ function evaluations, $s_{1}$ can successfully solve more problems than $s_{2}$.
When UPOQA converges on a given problem, the number of function evaluations for each element may differ. This discrepancy may arise from variations in the number of interpolation points required to initialize the surrogate models or from whether geometry improvement steps are performed on each element model during iterations. Let $t_{p,s}^{(i)}$ denote the number of function evaluations required for element $f_{i}$ when solver $s$ converges on problem $p$, $t_{p,s}^{\text{wst}}$ the maximum among $t_{p,s}^{(1)},\ldots,t_{p,s}^{(q)}$, and $t_{p,s}^{\text{avg}}$ their average. If the cost of evaluating the objective function is dominated by a specific element $f_{i_{*}}$, the overall cost of the algorithm will be determined by $t_{p,s}^{(i_{*})}$. Hence, in the most pessimistic scenario, the cost should be measured by $t_{p,s}^{\text{wst}}$ rather than $t_{p,s}^{\text{avg}}$. This is the meaning of the superscript wst (worst-case). Unless otherwise specified, the numerical experiments in subsequent sections generally use $t_{p,s}^{\text{wst}}$ to quantify the computational cost of UPOQA. When no ambiguity arises, $t_{p,s}^{\text{wst}}$ may be abbreviated as $t_{p}^{\text{wst}}$.
Since UPOQA leveraging partially separable structures is often significantly faster than algorithms that do not utilize such structures—potentially rendering comparisons between the two less meaningful—this paper introduces an alternative metric called the speed-up profile, defined as
$\mathrm{su}(\alpha)\triangleq\begin{cases}\frac{1}{\operatorname{card}(\mathcal{P})}\operatorname{card}\left(\left\{p\in\mathcal{P}:1\leq c_{p}\leq\alpha\land\left(t_{p}^{\text{single}}<\infty\lor t_{p}^{\text{wst}}<\infty\right)\right\}\right),&\quad\text{for}~{}\alpha\geq 1,\\ \frac{1}{\operatorname{card}(\mathcal{P})}\operatorname{card}\left(\left\{p\in\mathcal{P}:\alpha\leq c_{p}<1\land\left(t_{p}^{\text{single}}<\infty\lor t_{p}^{\text{wst}}<\infty\right)\right\}\right),&\quad\text{for}~{}0\leq\alpha<1,\end{cases}$
where
(15) $c_{p}\triangleq\frac{t_{p}^{\text{single}}/t_{p}^{\text{wst}}}{n_{p}~{}/\max\limits_{i=1,\ldots,q_{p}}n_{p,i}}$
and $t_{p}^{\text{single}}$ represents the number of function evaluations needed when the objective is treated as a single element and passed to UPOQA. Here, $n_{p,i}$ is the dimension of the element $f_{i}$ for problem $p$. Thus, $t_{p}^{\text{single}}/t_{p}^{\text{wst}}$ measures the actual speed-up achieved by exploiting the partially separable structure, while $n_{p}/\max\limits_{i=1,\ldots,q_{p}}n_{p,i}$ serves as a predicted speed-up ratio. To illustrate, consider an $n$-dimensional objective function formed by summing $n/\hat{n}$ low-dimensional subfunctions, each depending on distinct $\hat{n}$ variables. The Hessian is block-diagonal with identical $\hat{n}\times\hat{n}$ blocks. For gradient-based methods using forward-difference numerical gradients, exploiting this structure reduces per-iteration function evaluations from $O(n)$ to $O(\hat{n})$—each subgradient requires $\hat{n}+1$ evaluations rather than $n+1$ for the full gradient. Assuming comparable iteration counts, this induces a predicted speed-up ratio of $n/\hat{n}$. For general problems, the elemental dimensions $n_{1},\ldots,n_{q}$ may vary. In such cases, we assume that the element with largest dimension dominates the cost, making $n~{}/\max\limits_{i=1,\ldots,q_{p}}n_{i}$ a conservative prediction of the speed-up ratio. Notably, if any element has the same dimension as the objective function (i.e., $\max\limits_{i=1,\ldots,q_{p}}n_{i}=n$), UPOQA may fail to exploit any intrinsic low-dimensionality, resulting in a predicted speed-up ratio close to 1.
The expression (15) can be interpreted as a relative speed-up ratio, indicating whether UPOQA delivers the expected performance gain. A relative speed-up ratio greater than 1 suggests that revealing the partially separable structure not only allows the algorithm to exploit intrinsic low-dimensionality for reduced modeling cost but also enables the distillation of the primary nonconvexity into relatively low-dimensional elements, further lowering the actual cost. If UPOQA fails to converge regardless of whether the partially separable structure is utilized, such cases are excluded when computing $\mathrm{su}(\alpha)$.
To present the results more intuitively, we design $\mathrm{su}(\alpha)$ to be computed across two distinct regimes: $\alpha\geq 1$ and $10\leq\alpha<$ 1. For $\alpha\geq 1$, $\mathrm{su}(\alpha)$ represents the proportion of problems where UPOQA, after exploiting the partially separable structure, achieves a relative speed-up ratio $c_{p}\geq 1$ while maintaining $c_{p}\leq\alpha$. A higher value in this regime is desirable, and the corresponding curve in the speed-up profile plot should ideally cluster toward the upper-left region. Specifically, $\mathrm{su}(+\infty)$ quantifies the proportion of problems with $c_{p}\geq 1$. For the regime $10\leq\alpha<$ 1, $\mathrm{su}(\alpha)$ measures the proportion of problems where $c_{p}<1$ yet UPOQA attains a relative speed-up ratio no less than $\alpha$, while $\mathrm{su}(0)$ indicates the proportion of problems with $c_{p}<1$.
Furthermore, we always expect the plotted speed-up profile curve to exhibit a roughly balanced distribution across $\alpha=1$, and to rise rapidly on both sides of it. This implies that for the majority of problems, the predicted speed-up ratio $n_{p}/\max\limits_{i=1,\ldots,q_{p}}n_{p,i}$ approximates the actual speed-up ratio achieved.
3.2. Test Results on CUTEst Problems
3.2.1. Test Problems
Since its release in 1995, the CUTEst testing environment and problem set (Gould et al., 2015) has been widely used by researchers in numerical optimization for designing, testing, and comparing both constrained and unconstrained optimization algorithms. Each test problem is described in the standard input format (SIF), which utilizes the group partially separable structure (Conn et al., 2010) defined as:
(16) ${f(x)}=\sum_{i\in\mathcal{G}_{\text{obj }}}\frac{F_{i}\left(a_{i}(x),\omega_{i}\right)}{\sigma_{i}}+\frac{1}{2}x^{\top}Hx,$
where each term $F_{i}\left(a_{i}(x),\omega_{i}\right)/\sigma_{i}$ is called an objective-function group, $H\in{\mathbb{R}}^{n\times n}$ is symmetric, $\mathcal{G}_{\text{obj}}$ is an index set, and $F_{i}$ is a linear or nonlinear function of $a_{i}(x)$ and $\omega_{i}$. We employ the S2MPJ tool (Gratton and Toint, 2024) as the parser for SIF files to create partially separable versions of these problems with Python based on the group partially separable structure (16). In our implementation, each $F_{i}\left(a_{i}(x),\omega_{i}\right)/\sigma_{i}$ forms an element, and the quadratic term $x^{\top}Hx/2$ constitutes a separate element, together forming the partially separable structure of a problem.
We conducted a screening of available CUTEst problems, excluding those that are inconvenient for testing or have undesirable structures, including:
  • Constrained problems;
  • Problems with excessively large average elemental dimension $\overline{n}=\left(\sum_{i=1}^{q}n_{i}\right)/q$, indicating a lack of intrinsic low-dimensionality;
  • Problems with too many ($q\gg n$) or too few ($q=1$ or $2$) elements;
  • Problems requiring excessively long construction time;
  • Problems that are highly difficult to optimize, where none of the solvers in $\mathcal{S}$ can converge within the given evaluation budget;
  • Problems with excessively small or large dimensions.
Subsequently, we selected a test set of medium-scale problems, whose basic information is provided in Table 2 in Appendix A. This set comprises 85 partially separable problems with dimensions $n\in[21,~{}100]$, most of which are 50-dimensional. These problems typically contain a substantial number of elements, with many satisfying $q\geq n$. Whether this structure is sufficiently representative of real-world applications remains debatable. In contrast, the quantum variational problems discussed later satisfy $q\ll n$.
3.2.2. Test Results
We tested the UPOQA, UPOQA (single), BOBYQA, and L-BFGS-B (ffd) algorithms on the 85 problems included in Table 2 in Appendix A. Here, UPOQA (single) denotes optimizing the objective function as a whole element without exploiting the partially separable structure. The BOBYQA (Powell, 2009) implementation was from version 2.8.0 of Python’s nlopt library, while L-BFGS-B was from version 1.14.1 of the scipy library, using forward-finite-difference gradients for the objective. For each problem, the maximum number of function evaluations allowed was set to $\max\{1000n,10000\}$, with the same limit applied to each element function evaluation in UPOQA. For the convergence criterion (14), we selected $\varepsilon=10^{-1},~{}10^{-3},~{}10^{-5}$, and $10^{-7}$, obtaining the performance profiles in Figure 4d and the data profiles shown in Figure 5. Table 1 lists the number of function evaluations required for convergence on the five problems of largest dimensions in the test set, where $\infty$ indicates failure to converge within the evaluation budget.
Table 1. Function evaluations required for convergence by UPOQA, UPOQA (single), BOBYQA, and L-BFGS-B (ffd) on the five largest-dimensional problems from the test set in Table 2 in Appendix A, with convergence tolerances $\varepsilon=10^{-1},~{}10^{-3},~{}10^{-5}$, and $10^{-7}$. For UPOQA, the reported evaluations are the maximum evaluations across all elements. Here, $n$ denotes the problem dimension, $q$ denotes the number of elements in its partially separable structure, $n_{1},\ldots,n_{q}$ are dimensionalities of elements, and $\infty$ indicates failure to converge within the evaluation budget.
  Problem $n$ $q$ $\max n_{i}$ $\sum n_{i}/q$ $\log_{10}\varepsilon~{}$ Evaluations
10
UPOQA UPOQA (single) BOBYQA L-BFGS-B (ffd)
  HIMMELBI 100 20 5 5.00 -1 48 68 57 113
-3 83 97 72 113
-5 885 2357 $\infty$ 841
-7 20237 $\infty$ $\infty$ $\infty$
JANNSON3 100 102 2 1.01 -1 12 113 116 154
-3 47 429 976 358
-5 83 1309 2349 562
-7 108 1850 3640 1174
LUKSAN21LS 100 100 3 2.98 -1 11 246 115 103
-3 27 1033 588 307
-5 39 1318 838 511
-7 51 1517 1098 715
LINVERSE 99 147 4 2.98 -1 51 488 489 562
-3 139 1607 1815 1735
-5 145 2880 2486 2551
-7 152 3386 3712 3622
HYDC20LS 99 99 14 7.47 -1 11 101 101 52
-3 19 101 101 52
-5 19 101 101 256
-7 27 101 101 358
(a) $\varepsilon=10^{-1}$
(b) $\varepsilon=10^{-3}$
(c) $\varepsilon=10^{-5}$
(d) $\varepsilon=10^{-7}$
Figure 4. Performance profiles of UPOQA, UPOQA (single), BOBYQA, and L-BFGS-B (ffd) on the test problems from Table 2 in Appendix A, with convergence tolerances $\varepsilon=10^{-1},~{}10^{-3},~{}10^{-5}$, and $10^{-7}$.
[…]
Fully described in the text.
Figures 4d and 5, along with Table 1, demonstrate that UPOQA achieves significant acceleration by exploiting the partially separable structure. Across all convergence tolerances, it solves more problems with fewer evaluations. While UPOQA (single) solves slightly fewer problems than BOBYQA and L-BFGS-B (ffd), this is understandable since it was not specifically designed for single-objective optimization. For all tolerances, UPOQA solves at least 76.4% of problems with fewest evaluations, reaching 90.6% at $\varepsilon=10^{-1}$, indicating superior efficiency for low-accuracy solutions. Remarkably, UPOQA’s success rate declines only marginally as $\varepsilon$ decreases. At $\varepsilon=10^{-7}$, it solves approximately 5% more problems than the next best performer L-BFGS-B (ffd), while their performance was nearly identical at $\varepsilon=10^{-1}$. This demonstrates UPOQA’s robustness in high-accuracy scenarios.
(a) $\varepsilon=10^{-1}$
(b) $\varepsilon=10^{-3}$
(c) $\varepsilon=10^{-5}$
(d) $\varepsilon=10^{-7}$
Figure 5. Data profiles of UPOQA, UPOQA (single), BOBYQA, and L-BFGS-B (ffd) on the test problems from Table 2 in Appendix A, with convergence tolerances $\varepsilon=10^{-1},~{}10^{-3},~{}10^{-5}$, and $10^{-7}$.
[…]
Fully described in the text.
To further evaluate the acceleration effect of UPOQA, we present the speed-up profiles computed by UPOQA and UPOQA (single) in Figure 6. At a low convergence tolerance ($\varepsilon=10^{-1}$), the profile shows that among problems where at least one algorithm converged, 73.5% exhibited relative speed-up ratios between 0.5 and 2 (i.e., the gray-shaded region $[-1,1]$ on the horizontal axis), indicating reasonable alignment between predictions and empirical results. At the same time, UPOQA achieved lower speed-up than predicted on 57.6% of problems, outperformed predictions on 35.2%, while failing to converge within the budget on the remaining 7.2%—regardless of exploiting partially separable structures. Thus, UPOQA’s acceleration advantage is limited when only low-precision solutions are required. However, at higher precision ($\varepsilon=10^{-4}$), its acceleration becomes significantly enhanced, with a notably increased proportion of problems distributed on the $\log_{2}(\alpha)\geq 0$ side. When tolerance tightens further to $\varepsilon=10^{-7}$, the outperformance ratio shows no significant change, but the proportion of problems with $\log_{2}(\alpha)<0$ decreases markedly—a consequence of reduced solve success rates under stricter precision demands. Overall, the acceleration benefits of UPOQA are marginal in low-precision scenarios but become more pronounced in high-precision contexts.
Figure 6. Speed-up profiles computed by UPOQA and UPOQA (single) on the test problems from Table 2 in Appendix A, with convergence tolerances $\varepsilon=10^{-1},~{}10^{-4}$, and $10^{-7}$.
[…]
Fully described in the text.
3.3. Test Results on Quantum Variational Problems
With the advancement of quantum technology, the variational quantum algorithm (VQA) has garnered increasing attention as a hybrid quantum-classical approach for solving optimization problems (Cerezo et al., 2021). In brief, given an objective function, the first step of VQA involves constructing a parameterized quantum circuit on a quantum computer to represent this function. Classical derivative-free optimization algorithms are then employed to adjust these parameters and ultimately find the minimum of the objective. Now, let us consider a quantum variational problem:
(17) $\min_{\theta_{1},\ldots,\theta_{p}\ \in\ {\mathbb{R}}^{n_{\theta}}}\quad\sum_{i=1}^{p}w_{i}\left\langle\phi_{i}\left|U(\theta_{i})^{\dagger}\widehat{H}U(\theta_{i})\right|\phi_{i}\right\rangle+\lambda\sum_{i<j}\left|\langle\phi_{i}\left|U(\theta_{i})^{\dagger}U(\theta_{j})\right|\phi_{j}\rangle\right|^{2},$
where $\widehat{H}$ denotes the Hamiltonian operator of the molecular system, $U(\theta_{1}),\ldots,U(\theta_{p})$ represent the parameterized ansatz quantum circuits with parameters $\theta_{1},\ldots,\theta_{p}$ respectively, $\phi_{1},\ldots,\phi_{p}$ are the initial quantum states, $w_{1},\ldots,w_{p}>0$ are the weights, and $\lambda>0$ is the penalty coefficient. The goal of this problem is to determine the states corresponding to the smallest $p$ energy levels of a given molecular system.
In equation (17), each term under the summation represents the measurement outcome of a quantum circuit, which is typically quite costly. We treat each term in (17) as a black box and reformulate the problem as:
(18) $\min_{\theta_{1},\ \ldots,\ \theta_{p}\ \in\ {\mathbb{R}}^{n_{\theta}}}\ \sum_{i=1}^{p}f_{i}(\theta_{i})+\sum_{1\leq i<j\leq p}f_{ij}(\theta_{i},\theta_{j}),$
where $f_{i}(\theta_{i})=w_{i}\left\langle\phi_{i}\left|U(\theta_{i})^{\dagger}\widehat{H}U(\theta_{i})\right|\phi_{i}\right\rangle$, $f_{ij}(\theta_{i},\theta_{j})=\lambda\left|\left\langle\phi_{i}\left|U(\theta_{i})^{\dagger}U(\theta_{j})\right|\phi_{j}\right\rangle\right|^{2}$. Clearly, this is a coordinate partially separable problem, where the objective function is $np$-dimensional and consists of $p(p+1)/2$ elements with a maximum dimensionality of $2n$ and exhibits a dense Hessian matrix. Similar structures are common in other quantum variational problems (Bierman et al., 2022, 2023).
We conducts numerical experiments on problem (18) for $\text{H}_{2}$ and a model consisting of four hydrogen atoms arranged in a square lattice, which is denoted as $\text{H}_{4}$, in the STO-3G basis with UCCSD (Romero et al., 2018) ansatz circuits. The UCCSD block pattern is repeated for three times in order to increase the expressiveness of the ansatz. We set $p=3$, indicating that the goal is to compute the three lowest-energy excited states and their state energies, with weights set as $w_{i}=i$ for $i=1,2,3$. The set of initial states $\phi_{1},\phi_{2},\phi_{3}$ and all molecular Hamiltonians are chosen to be the Hartree-Fock state and low-lying single-particle excitations above it. The starting points $\theta_{0,1},\theta_{0,2},\theta_{0,3}$ are randomly sampled according to a uniform distribution on $[-2\pi,2\pi)$. Both experiments are performed in a noiseless environment. The experimental code is implemented in Python using qiskit version 0.37.0, with quantum circuit designs consistent with those in the quantum orbital minimization method (qOMM(Bierman et al., 2022). In the $\text{H}_{2}$ experiment, the interatomic distance is set to 0.735 Å, and the penalty coefficient is empirically chosen as $\lambda=12$. For the $\text{H}_{4}$ experiment, the interatomic distance is 1.23 Å, with $\lambda=30$.
(a) $\text{H}_{2}$ molecule
(b) $\text{H}_{4}$ molecule
Figure 7. Objective function value descent curves when optimizing the noiseless quantum variational problem (18) for $\text{H}_{2}$ and $\text{H}_{4}$ molecular systems using UPOQA, UPOQA (single), BOBYQA, and L-BFGS-B (ffd).
[…]
Fully described in the text.
The experimental results from both tests are shown in Fig 7b. UPOQA demonstrates superior performance on both problems, with more pronounced acceleration effects in the $\text{H}_{2}$ experiment. Since the predicted acceleration rate is $p/2=1.5$, the results align well with theoretical predictions. In both experiments, all solvers encountered ”plateau periods” where the objective function value stagnated, but UPOQA consistently escaped these phases faster and achieved solutions with fewer function evaluations. These results indicate that leveraging the internal structure of problem (18) indeed leads to acceleration, and suggest that UPOQA could deliver outstanding performance on larger-scale problems and other general quantum variational problems.
4. Conclusion
We propose the UPOQA algorithm for derivative-free optimization of partially separable problems. Based on quadratic interpolation models and a trust-region framework, UPOQA constructs underdetermined quadratic models for each element function, employing the Steinmetz projection and a modified projected gradient method to solve structured trust-region subproblems. The implementation incorporates Powell’s techniques (Powell, 2004b, 2008), resulting in low iteration costs and strong robustness. The trust-region radius management strategy is modified from Shahabuddin’s criterion (Shahabuddin, 1996, Section 2.4). Additionally, UPOQA features practical functionalities such as starting point search, restart mechanism, and hybrid black-white-box optimization. Comprehensive numerical experiments on CUTEst problems and a set of quantum variational problems validate the algorithm’s effectiveness and robustness in problem-solving.
We acknowledge several aspects of UPOQA that warrant further research and improvement. The modified projected gradient method currently exhibits limited single-step descent rates and requires excessive iterations, leading to significant runtime overhead in some scenarios. Theoretical analysis remains incomplete, particularly regarding convergence rates and their relationship with partially separable structures. Experimentally, we aspire to test the algorithm on more real-world applications, as the artificially constructed partially separable structures from CUTEst test problems may lack sufficient representativeness. Future enhancements could potentially incorporate state-of-the-art DFO techniques such as sample averaging and regression models (Cartis et al., 2019), as well as preconditioning and subspace methods (Zhang, 2025), which might further improve UPOQA’s performance.
Appendix A Test Problems
Table 2 lists all the test problems used in Section 3.2. These problems were extracted from the CUTEst problem set (Gould et al., 2015) using the S2MPJ tool (Gratton and Toint, 2024).
Table 2. CUTEst test problems used in Section 3.2, where $n$ denotes the problem dimension, $q$ denotes the number of elements, and $n_{1},\ldots,n_{q}$ are the dimensions of elements.
Problem $n$ $q$ $\max n_{i}$ $\sum n_{i}/q$ Problem $n$ $q$ $\max n_{i}$ $\sum n_{i}/q~{}$
ANTWERP 27 19 19 7.74 ARWHEAD 50 98 2 1.50
BDQRTIC 50 92 5 3.00 BIGGSB1 50 51 2 1.96
BROYDN3DLS 50 50 3 2.96 BROYDNBDLS 50 50 7 6.68
CHNROSNB 50 98 2 1.50 CHNRSNBM 50 98 2 1.50
COSINE 50 49 2 2.00 CURLY10 50 50 11 9.90
CURLY20 50 50 21 16.80 CURLY30 50 50 31 21.70
CVXQP1 50 50 3 2.94 CYCLOOCFLS 86 60 6 5.73
CYCLOOCTLS 90 60 6 6.00 DIXON3DQ 50 50 2 1.96
DQRTIC 50 50 1 1.00 DTOC2 88 15 6 5.87
EG2 50 50 2 1.96 ENGVAL1 50 98 2 1.50
ERRINROS 50 98 2 1.50 ERRINRSM 50 98 2 1.50
EXTROSNB 50 50 2 1.98 FLETBV3M 50 102 50 1.96
FLETCBV2 50 151 2 1.32 FLETCHCR 50 98 2 1.50
FREUROTH 50 98 2 2.00 GILBERT 50 50 1 1.00
HATFLDC 25 25 2 1.92 HIMMELBI 100 20 5 5.00
HUESmMOD 50 50 1 1.00 HYDC20LS 99 99 14 7.47
HYDCAR6LS 29 29 14 6.69 INDEFM 50 98 3 1.98
JANNSON3 100 102 2 1.01 JANNSON4 50 52 2 1.02
LEVYMONT10 50 100 2 1.49 LEVYMONT6 50 100 2 1.49
LIARWHD 50 100 2 1.49 LINVERSE 99 147 4 2.98
LUKSAN21LS 100 100 3 2.98 LUKVLI10 50 50 2 2.00
LUKVLI11 50 64 2 1.25 LUKVLI13 50 48 2 1.67
LUKVLI14 50 64 2 1.25 LUKVLI1 50 98 2 1.50
LUKVLI2 50 125 2 1.62 LUKVLI3 50 96 2 2.00
LUKVLI4C 50 120 2 1.60 LUKVLI5 52 50 3 3.00
LUKVLI8 50 40 5 4.00 LUKVLI9 50 51 50 2.45
METHANB8LS 31 31 11 6.29 METHANL8LS 31 31 11 6.29
MOREBV 50 50 3 2.96 NCB20B 50 50 20 12.78
NCB20 60 51 30 12.75 NONDQUAR 50 50 3 2.96
NONSCOMP 50 50 2 1.98 OSCIGRAD 50 50 3 2.96
OSCIPATH 50 50 2 1.98 PENALTY1 50 51 50 1.96
PENALTY2 50 100 50 1.98 QING 50 50 1 1.00
RAYBENDL 62 30 4 4.00 SANTALS 21 23 4 3.52
SBRYBND 50 50 7 6.68 SCHMVETT 50 48 3 3.00
SCOSINE 50 49 2 2.00 SINEALI 50 50 2 1.98
SINQUAD2 50 50 3 2.94 SINROSNB 50 50 2 1.98
SPARSINE 50 50 6 5.58 SPARSQUR 50 50 6 5.58
SSBRYBND 50 50 7 6.68 STRTCHDV 50 49 2 2.00
TOINTGOR 50 83 5 1.81 TOINTGSS 50 48 3 3.00
TOINTPSP 50 83 5 1.81 TOINTQOR 50 83 5 1.81
TQUARTIC 50 50 2 1.98 TRIDIA 50 50 2 1.98
VARDIM 50 52 50 2.88 VAREIGVL 51 51 50 13.88
YAO 52 52 1 1.00
Appendix B Performance of UPOQA Using Unstructured Trust Regions
We also compared the performance of UPOQA with and without structured trust regions. When structured trust regions are not employed, all the radii $\Delta_{k,1},\ldots,\Delta_{k,q}$ remain identical, and the trust-region subproblem (6) reduces to an optimization problem within a spherical region. We solve it using the truncated conjugate gradient method with a boundary improvement step, as implemented in COBYQA (Ragonneau, 2023). This variant of UPOQA is labeled UPOQA (non-struct).
(a) $\varepsilon=10^{-3}$
(b) $\varepsilon=10^{-7}$
Figure 8. Performance profiles of UPOQA and its unstructured-trust-region variant UPOQA (non-struct) on the test problems from Table 2 in Appendix A, with convergence tolerances $\varepsilon=10^{-3}$ and $10^{-7}$.
[…]
Fully described in the text.
We plot the performance profiles for both algorithms under convergence tolerances of $\varepsilon=10^{-3}$ and $10^{-7}$, as shown in Figure 8b. At $\varepsilon=10^{-3}$, UPOQA significantly outperforms UPOQA (non-struct). With structured trust regions, UPOQA achieves faster convergence on 65.8% of the problems, whereas UPOQA (non-struct) exhausts its evaluation budget without reaching the same solution accuracy as UPOQA on 11.7% of the problems. At $\varepsilon=10^{-7}$, the performance gap narrows slightly, with UPOQA converging first on 56.4% of the problems. These results demonstrate the effectiveness of replacing the unstructured trust region with a structured one.